# Bringing memory to Boolean networks: a unifying framework

### v1

@@ 2.1 Asynchronous updates
We begin with the simplest update mode, namely asynchronous updates. In this case, a component is updated at each step and the result is calculated and applied to the last configuration visited. Then, we select $i^{a}$ and apply $f_{i^{a}}$ to the current configuration $x^{a-1}$. In other words, $x\to_{\mathtt{A}}y$ if and only if $y\in N^{out}(\mathtt{A}(f);x)$. Thus, the memory consists only of the current configuration. In our framework, the definition is as follows.
@@
$\mathsf{A}_{1}(x)$ =$=\neg x_{1}\lor\neg x_{2},$
$\mathsf{A}_{2}(x)$ =$=\neg x_{1}\lor\neg x_{2}.$
From the asynchronous graph of $\mathsf{A}$ (shown below), we can see that $1100\to_{\mathtt{A}}^{*}$ 11. A possible asynchronous trajectory $0100\to_{\mathtt{A}}10\to_{\mathtt{A}}11\to_{\mathtt{A}}$ 01 is detailed in the table below.
00000000010101011010101011111111
@@ 00 $00$
$1$ $1$ $00$ $00$ $10$
$2$ $2$ $10$ $10$ $11$
$3$ $1$ $11$ $11$ $01$
2.2 History-based updates
@@
The asynchronous graph of $\mathsf{H}$ is given below, as well as a possible $\mathtt{H}$-trajectory. The table is a possible representation of the trajectory $101000\to_{\mathtt{H}}100\to_{\mathtt{H}}110\to_{\mathtt{H}}111\to_{\mathtt{H}}$ 101. It shows how $i^{a}$, $s^{a}$ and $t^{a}$ are chosen in each step. A new equivalent representation is proposed in Figure 2.
000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
@@ 00 $000$
$1$ $1$ $000$ $000$ $100$
$2$ $2$ $100$ $100$ $110$
$3$ $3$ $110$ $110$ $111$
$4$ $2$ $000$ $111$ $101$
2.3 Trapping updates
So far, all our updates applied $f_{i^{a}}$ to some source configuration $s^{a}$ and copied the rest of the current configuration: $x^{a}_{-i^{a}}=x^{a-1}_{-i^{a}}$. Let us now consider an update mode in which the rest of the configuration can be taken from the history of the trajectory. Trapping updates are the natural generalisation of history-based updates, where both the source $s^{a}$ and the target $t^{a}$ are chosen from the set $\{x^{0},\dots,x^{a-1}\}$. The term “Trapping updates” come from the fact, proved in the sequel, that $x\to_{\mathtt{T}}^{*}y$ if and only if $y\in T_{f}(x)$. In our framework, this turns out to be formalised as follows.
@@ 2.3 Trapping updates
  • target $t^{a}\in\{x^{0},\dots,x^{a-1}\}$;
  • new configuration $x^{a}=(f_{i^{a}}(s^{a}),t^{a}_{-i^{a}})$.
In trapping updates, it is possible to account for possible delays in updating a component (due to the fact that $s^{a}\in\{x^{0},\dots,x^{a-1}\}$), and it is also possible that the result of the update is applied on a past configuration (i.e. $t^{a}\in\{x^{0},\dots,x^{a-1}\}$), becoming the new current configuration.
@@
The asynchronous graph of $\mathsf{T}$ is given below, as well as a possible $\mathtt{T}$-trajectory showing how $0100\to^{*}_{\mathtt{T}}$ 01. Note how $x^{3}$ is computed with $t^{3}=00\neq x^{2}$.
00000000010101011010101011111111
@@ 00 $00$
$1$ $1$ $00$ $00$ $10$
$2$ $2$ $10$ $10$ $11$
$3$ $2$ $11$ $00$ $01$
2.4 Most permissive updates
In history-based updates, we assumed that the delays to transmit the information from any $j$ to $i^{a}$ was the same for all $j$. We can generalise this by allowing a different delay for any $j$, hence $f_{i^{a}}$ is now applied to some $s^{a}$ where each $s^{a}_{j}$ can be chosen from $\{x^{0}_{j},\dots,x^{a-1}_{j}\}$; this is equivalent to $s^{a}\in[x^{0},\dots,x^{a-1}]$. We now obtain the so-called most permissive updates introduced in [14], which in our framework becomes:
@@ 2.4 Most permissive updates
  • target $t^{a}=x^{a-1}$;
  • new configuration $x^{a}=(f_{i^{a}}(s^{a}),x_{-i^{a}}^{a-1})$.
In [14], most permissive updates were motivated by the abstraction of quantitative systems by Boolean networks: the fact that the source configuration can be chosen within the subcube $[x^{0},\dots,x^{a-1}]$ aims at capturing the potential heterogeneity of influence thresholds and time scales between component updates. They demonstrated that most permissive updates allow capturing all the dynamics possible in any quantitative refinement of the BN, while the (general) asynchronous mode hinder the prediction of several observed trajectories.
@@
$\mathsf{M}(x)$ $=x\text{ otherwise.}$.
The asynchronous graph of $\mathsf{M}$ is given below, as well as a possible $\mathtt{M}$-trajectory showing how $111000\to^{*}_{\mathtt{M}}$ 111.
000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
@@ 00 $000$
$1$ $1$ $000$ $000$ $100$
$2$ $2$ $100$ $100$ $110$
$3$ $3$ $010$ $110$ $111$
2.5 Subcube-based updates
Combining trapping and most permissive updates, subcube-based updates are defined on the basis of the subcube containing the configurations already visited: both the source $s^{a}$ and the target $t^{a}$ are chosen from the subcube $[x^{0},\dots,x^{a-1}]$.
@@ 2.5 Subcube-based updates
The asynchronous graph of $\mathsf{S}$ is given below, as well as a possible $\mathtt{S}$-trajectory showing how $001000\to^{*}_{\mathtt{S}}$ 001.
000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
@@ 00 $000$
$1$ $1$ $000$ $000$ $100$
$2$ $2$ $100$ $000$ $010$
$3$ $3$ $110$ $000$ $001$
2.6 Interval updates
The interval update was initially introduced for Petri nets in [4], and later adapted to BNs in [5], as a mean to capture transitions that could occur while a component is changing of state, i.e., while the component is committed to change, but the change has not been applied yet. One way to model this mode is to consider that each component $i\in V$ is decoupled in two nodes: a write node storing the next value and a read node for the current value. Whenever a state change for a component is triggered, only its write node is updated first. Then, the read node will be updated later, but in the meantime, other components can change of value based on the read node state.
In our framework, the definition of the interval update requires coupling each configuration $x^{a}$ of a trajectory with a vector $V^{a}$ associating to each component the time to read its state. Intuitively, for any component $j$, having $V^{a}_{j}<a-1$ enables to use its state prior to its latest state change (thus, $j$ state change is pending), whereas $V^{a}_{j}=a-1$ ensures that its state change has been completed, and other components read it. When a component $i$ is updated at time $a$, we impose that $V^{a}_{i}=a-1$, i.e., the most recent state of $i$ must be used. This captures the fact that a component has to have completed its state change before being updated again, which is enforced by the original definition of the Interval update mode [4, 5]. We also enforce $V^{a}\geq V^{a-1}$, which forbids the scenario described above for history-based updates.
Interval updates $\mathtt{I}$:
  • vector $V^{a}\in\mathbb{R}^{n}$ such that $V^{0}=0^{n}$, $V^{a}\geq V^{a-1}$, $V^{a}\leq(a-1)^{n}$ and $V^{a}_{i^{a}}=a-1$;
  • source $s^{a}$ with $s^{a}_{j}=x^{V^{a}_{j}}_{j}$ for all $j\in[n]$;
  • target $t^{a}=x^{a-1}$;
  • new configuration $x^{a}=(f_{i^{a}}(s^{a}),x^{a-1}_{-i^{a}})$.
Notice that, in general, several $V^{a}$ can be possible, modelling different schedule of state change completion, possibly enabling reaching different configurations. The asynchronous update mode, where no memory of previous states is used, corresponds to the case whenever $V^{a}=(a-1)^{n}$ for any time $a\geq 1$.
@@ 2.6 Interval updates
The asynchronous graph of $\mathsf{I}$ is shown below. We also show an example of $\mathtt{I}$-trajectory.
000000000000001001001001010010010010011011011011100100100100101101101101110110110110111111111111
@@ 00 $\begin{bmatrix}0&0&0\end{bmatrix}$ $000$
$1$ $1$ $\begin{bmatrix}0&0&0\end{bmatrix}$ $000$ $000$ $100$
$2$ $3$ $\begin{bmatrix}0&0&1\end{bmatrix}$ $000$ $100$ $101$
$3$ $2$ $\begin{bmatrix}0&2&1\end{bmatrix}$ $000$ $101$ $111$
$4$ $1$ $\begin{bmatrix}3&3&1\end{bmatrix}$ $110$ $111$ $011$
In the trajectory shown, all components are first updated once with using the memory of the initial state. Then, the first component is updated again based on the most recent value of components 1, 2, but using the previous state of component 3 ($V^{4}_{3}=1$, i.e., the time just before the last update of $3$). Indeed, with $V^{4}=\begin{bmatrix}3&3&1\end{bmatrix}$, we obtain $s^{a}=x_{1}^{3}x_{2}^{3}x_{3}^{1}=110$ and $\mathsf{I}(110)_{1}=0$. Thus, $x^{4}=011$.
2.7 Cuttable updates
Finally, we consider the cuttable update mode introduced in [13], which generalizes the interval update mode by considering a finer description of the delay of state changes. Essentially, with the interval update mode, a component can have a delay to apply (transmit) its state change. Once it is transmitted (the read node is updated), all the components of the network will read its new state. The cuttable update mode allows to have different delays of transmission for each component of the network. Thus, instead of having a read node per component, there is a read node per pair of components. The read node copies the state of the (unique) write node of components asynchronously. Hence, after a change of state of a component $i$ has been fired, there can be moments whenever a component $j$ has access to its newer state, while another component $k$ has access to its previous state.
Cuttable updates $\mathtt{C}$:
  • matrix $C^{a}\in\mathbb{R}^{n\times n}$ such that $C^{0}=0$, $C^{a}\geq C^{a-1}$ and, for all $i,j\in[n]$, $10\leq C^{a}_{i,j}\leq a-$ 1;
  • source $s^{a}$ with $s^{a}_{j}=x^{C^{a}_{i^{a},j}}_{j}$ for all $j\in[n]$;
  • target $t^{a}=x^{a-1}$;
  • new configuration $x^{a}=(f_{i^{a}}(s^{a}),x^{a-1}_{-i^{a}})$.

### v2

@@ 2.2.1 Asynchronous updates
In this case, a component is updated at each step and the result is calculated and applied to the last configuration visited. Then, we select $i^{a}$ and apply $f_{i^{a}}$ to the current configuration $x^{a-1}$. In other words, $x\to_{\mathtt{A}}y$ if and only if $(x,y)$ is an arc in the asynchronous graph of $f$. Thus, the memory consists only of the current configuration. In our framework, the definition is as follows.
@@
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
@@
$\mathsf{A}_{1}(x)$ $=\neg x_{1}\lor\neg x_{2},$
$\mathsf{A}_{2}(x)$ $=\neg x_{1}\lor\neg x_{2}.$
From the asynchronous graph of $\mathsf{A}$ (shown below), we can see that $00\to_{\mathtt{A}}^{*}11$. A possible asynchronous trajectory $00\to_{\mathtt{A}}10\to_{\mathtt{A}}11\to_{\mathtt{A}}01$ is detailed in the table below.
0000010110101111
@@ 0 $00$
11 11 $00$ $00$ $10$
22 22 $10$ $10$ $11$
33 11 $11$ $11$ $01$
2.2.2 History-based updates
@@
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
@@
The asynchronous graph of $\mathsf{H}$ is given below, as well as a possible $\mathtt{H}$-trajectory. The table is a possible representation of the trajectory $000\to_{\mathtt{H}}100\to_{\mathtt{H}}110\to_{\mathtt{H}}111\to_{\mathtt{H}}101$. It shows how $i^{a}$, $s^{a}$ and $t^{a}$ are chosen in each step. A new equivalent representation is proposed in Figure 1.
000000001001010010011011100100101101110110111111
@@ 0 $000$
11 11 $000$ $000$ $100$
22 22 $100$ $100$ $110$
33 33 $110$ $110$ $111$
44 22 $000$ $111$ $101$
2.2.3 Trapping updates
So far, all our updates applied $f_{i^{a}}$ to some source configuration $s^{a}$ and copied the rest of the current configuration: $x^{a}_{-i^{a}}=x^{a-1}_{-i^{a}}$. Let us now consider an update mode in which the rest of the configuration can be taken from the history of the trajectory. Trapping updates are the natural generalisation of history-based updates, where both the source $s^{a}$ and the target $t^{a}$ are chosen from the set $\{x^{0},\dots,x^{a-1}\}$. Thus, it allows accounting for possible delays in updating a component, and it allows the result of the update to be applied on a past configuration, becoming the new current configuration. The term “Trapping updates” come from the fact, proved in the sequel, that $x\to_{\mathtt{T}}^{*}y$ if and only if $y\in T_{f}(x)$. In our framework, this turns out to be formalised as follows.
@@ 2.2.3 Trapping updates
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
@@
  • target $t^{a}\in\{x^{0},\dots,x^{a-1}\}$.
@@
The asynchronous graph of $\mathsf{T}$ is given below, as well as a possible $\mathtt{T}$-trajectory showing how $00\to^{*}_{\mathtt{T}}01$. Note how $x^{3}$ is computed with $t^{3}=00\neq x^{2}$.
0000010110101111
@@ 0 $00$
11 11 $00$ $00$ $10$
22 22 $10$ $10$ $11$
33 22 $11$ $00$ $01$
2.2.4 Most permissive updates
We now focus on the most permissive updates introduced in [13]. This updating technique can be defined using dynamic states or hypercubes, let us recall the definition in this first form.
In this update mode, the change of state of the components is broken down into two stages, with a dynamic state indicating an increase $\nearrow$ or decrease $\searrow$. The set of possible states of a component is denoted $\mathbb{P}=\{0,\nearrow,\searrow,1\}$. When a component is in a dynamic state, the other components in the network can arbitrarily read 0 or $1$. The function $\gamma:\mathbb{P}^{n}\to 2^{\mathbb{B}^{n}}$ associates with any configuration $x\in\mathbb{P}^{n}$ the set of configurations that correspond to it. Then, $\gamma(x)=\{x^{\prime}\in\mathbb{B}^{n}\mid\forall i\in[n],x_{i}\in\mathbb{B}\Rightarrow x^{\prime}_{i}=x_{i}\}$. In this semantics, it is therefore defined that, $\forall x,y\in\mathbb{P}^{n}$, $x\to y$ iff $\exists i\in[n]:\Delta(x,y)=\{i\}$ and
[…]
The most permissive update mode considers that a component in the process of changing of state can be viewed as in superposition of both states. As such, in our framework, this is equivalent to allowing a different delay for any $j$, hence $f_{i^{a}}$ is now applied to some $s^{a}$ where each $s^{a}_{j}$ can be chosen from $\{x^{0}_{j},\dots,x^{a-1}_{j}\}$; this is equivalent to $s^{a}\in[x^{0},\dots,x^{a-1}]$.
The most permissive update mode can therefore be formalised in our framework as follows.
@@ 2.2.4 Most permissive updates
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
@@
  • target $t^{a}=x^{a-1}$.
The correspondence between the two definitions is given in Sect. A.1.
In [13], most permissive updates were motivated by the abstraction of quantitative systems by Boolean networks: the fact that the source configuration can be chosen within the subcube $x^{0},\dots,x^{a-1}]$ aims at capturing the potential heterogeneity of influence thresholds and time scales between component updates. They demonstrated that most permissive updates allow capturing all the dynamics possible in any quantitative refinement of the BN, while the (general) asynchronous mode hinder the prediction of several observed trajectories.
@@
$\mathsf{M}(x)$ $=x\text{ otherwise.}$
The asynchronous graph of $\mathsf{M}$ is given below, as well as a possible $\mathtt{M}$-trajectory showing how $000\to^{*}_{\mathtt{M}}111$.
000000001001010010011011100100101101110110111111
@@ 0 $000$
11 11 $000$ $000$ $100$
22 22 $000$ $100$ $110$
33 33 $010$ $110$ $111$
2.2.5 Subcube-based updates
Subcube-based updates are defined similarly to trapping updates, but instead consider the subcube $x^{0},\dots,x^{a-1}]$ containing the configuration already visited for both the source $s^{a}$ and the target $t^{a}$.
@@ 2.2.5 Subcube-based updates
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
@@
The asynchronous graph of $\mathsf{S}$ is given below, as well as a possible $\mathtt{S}$-trajectory showing how $000\to^{*}_{\mathtt{S}}001$.
000000001001010010011011100100101101110110111111
@@ 0 $000$
11 11 $000$ $000$ $100$
22 22 $100$ $000$ $010$
33 33 $110$ $000$ $001$
2.2.6 Interval updates
The interval update was initially introduced for Petri nets in [3], and later adapted to BNs in [4], as a mean to capture transitions that could occur while a component is changing of state, i.e., while the component is committed to change, but the change has not been applied yet. This was formalised by proposing to consider that each component $i\in[n]$ is decoupled in two nodes: a write node (2i1)(2i-1) storing the next value and a read node $(2i)$ for the current value. Whenever a state change for a component is triggered, only its write node is updated first. Then, the read node will be updated later, but in the meantime, other components can change of value based on the read node state.
Let us define a function $\tau:\mathbb{B}^{2n}\to\mathbb{B}^{n}$ which map a configuration of the interval semantics to a configuration of the BN $f$ by projecting on the read nodes, i.e., $\tau(z)_{i}=z_{2i}$ for every $i\in[n]$. Given a BN $f$ with $n$ components, $\tilde{f}$ is a BN with $2n$ components where $\forall i\in[n]$, we have $\tilde{f}_{2i}(z)=z_{2i-1}$ and $\tilde{f}_{2i-1}(z)=\begin{cases}f_{i}(\tau(z))&\text{if }z_{2i}=z_{2i-1}\\ z_{2i-1}&\text{otherwise}\end{cases}$.
@@ 2.2.6 Interval updates
We can obtain the following possible sequence of asynchronous iterations of $\tilde{f}$ with the interval semantics. We insert spaces in the configurations to facilitate the identification of the two read/write components corresponding to each of the original components.
[…]
Hence, the configuration $011$ of $\mathsf{I}$ is reachable from $000$ according to this update mode.
In our framework, the definition of the interval update requires coupling each configuration $x^{a}$ of a trajectory with a vector $V^{a}$ associating to each component the time to read its state. Intuitively, for any component $j$, having $V^{a}_{j}<a-1$ enables to use its state prior to its latest state change (thus, $j$ state change is pending), whereas $V^{a}_{j}=a-1$ ensures that its state change has been completed, and other components read it. When a component $i$ is updated at time $a$, we impose that $V^{a}_{i}=a-1$, i.e., the most recent state of $i$ must be used. This captures the fact that a component has to have completed its state change before being updated again, which is enforced by the original definition of the Interval update mode [3, 4]. We also enforce $V^{a}\geq V^{a-1}$ (i.e., for any component $j$, $V^{a}_{j}\geq V^{a-1}_{j}$), which forbids the scenario described above for history-based updates.
Interval updates $\mathtt{I}$:
All trajectories $(x^{0},\dots,x^{l})$ obtained by $x^{0}\in\mathbb{B}^{n}$ and $w^{a}=(i^{a},s^{a},t^{a})\in[n]\times\mathbb{B}^{n}\times\mathbb{B}^{n}$ for $1\leq a\leq l$, where
  • vector $V^{a}\in\mathbb{N}^{n}$ such that $V^{0}=0^{n}$, $V^{a}\geq V^{a-1}$, $V^{a}\leq(a-1)^{n}$ and $V^{a}_{i^{a}}=a-1$;
  • source $s^{a}$ with $s^{a}_{j}=x^{V^{a}_{j}}_{j}$ for all $j\in[n]$;
  • target $t^{a}=x^{a-1}$.
Notice that, in general, several $V^{a}$ can be possible, modelling different schedule of state change completions, possibly enabling reaching different configurations. The asynchronous update mode, where no memory of previous states is used, corresponds to the case whenever $V^{a}=(a-1)^{n}$ for any time $a\geq 1$. The correspondence between the two definitions is given in Sect. A.2. Intuitively, the memory-based trajectories correspond to the projection of sequences of interval iterations over the write nodes and $V^{a}$ reflect the read nodes.
Example 2.7.
Let us consider the network of the previous example. The asynchronous graph of $\mathsf{I}$ is shown below. We also show an example of $\mathtt{I}$-trajectory.
000000001001010010011011100100101101110110111111
@@ 0 $\begin{bmatrix}0&0&0\end{bmatrix}$ $000$
11 11 $\begin{bmatrix}0&0&0\end{bmatrix}$ $000$ $000$ $100$
22 33 $\begin{bmatrix}0&0&1\end{bmatrix}$ $000$ $100$ $101$
33 22 $\begin{bmatrix}0&2&1\end{bmatrix}$ $000$ $101$ $111$
44 11 $\begin{bmatrix}3&3&1\end{bmatrix}$ $110$ $111$ $011$
In the trajectory shown, all components are first updated once with using the memory of the initial state (as done in the previous example). Then, the first component is updated again based on the most recent value of components 1, 2, but using the previous state of component 3 ($V^{4}_{3}=1$, i.e., the time just before the last update of $3$). Indeed, with $V^{4}=\begin{bmatrix}3&3&1\end{bmatrix}$, we obtain $s^{4}=x_{1}^{3}x_{2}^{3}x_{3}^{1}=110$ and $\mathsf{I}(110)_{1}=0$. Thus, $x^{4}=011$. It can be seen that the trajectories in this example and the previous one are equivalent. Intuitively, updating the read component based on the write component corresponds to updating the value in $V^{a}$.
2.2.7 Cuttable updates
Finally, we consider the cuttable update mode introduced in [12], which generalizes the interval update mode by considering a finer description of the delay of state changes. Essentially, with the interval update mode, a component can have a delay to apply (transmit) its state change. Once it is transmitted (the read node is updated), all the components of the network will read its new state. The cuttable update mode allows having different delays of transmission for each component of the network. Thus, instead of having a read node per component, there is a read node per pair of components. The read node copies the state of the (unique) write node of components asynchronously. Hence, after a change of state of a component $i$ has been fired, there can be moments whenever a component $j$ has access to its newer state, while another component $k$ has access to its previous state.
Let us recall the definition of an $E$-extension of a Boolean network $f$. For simplicity, let us define $E$ as the set of pairs $(i,j)$ such that component $i$ influences the value of component $j$ (i.e., $x_{i}$ is involved in the function $f_{j}$). It is possible to define the BN $\tilde{f}:\mathbb{B}^{n+|E|}\to\mathbb{B}^{n+|E|}$ such that
$\tilde{f}_{i}(x)$ $=f_{i}(\tau^{i}(x))$  $\quad\text{ for all }\ i\in[n],$
$\tilde{f}_{e}(x)$ $=x_{j}$  $\quad\text{ if }e=(j,k)\in E,$
where $\tau^{i}:\mathbb{B}^{n+|E|}\to\mathbb{B}^{n}$ is defined, for all $j\in[n]$ as:
$\tau^{i}(x)_{j}=\begin{cases}x_{(j,i)}&\text{ if }(j,i)\in E\\ x_{j}&\text{ otherwise. }\end{cases}$